# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v1

@@ 4.2 Oracle complexity with a two-sided oracle
We now turn to discuss the proof of Thm. 4.2. We begin by noting that the proof technique of Thm. 4.1 is of no use here, as it is based on revealing the rows of $A$ to the algorithm one at a time. This is not possible with a two-sided oracle, which allows us (for instance) to compute an arbitrary weighted combination of all rows of $A$ using a single query. A second difficulty is that we consider a very simple class of homogeneous bilinear functions, which means that any lower bound necessarily has to be of this form (as opposed to more general min-max optimization problems, where for lower bounds we can use functions with a richer structure). To handle these difficulties, we introduce a different proof technique, which still forces the algorithm to discover information about $A$ in an incremental manner, but in terms of subspaces rather than rows. Specifically, we will utilize the following randomized construction of $A$:
@@
    • Let $\mathbf{u}_{t}$ be an arbitrary unit vector in $\mathbb{R}^{d}$, which is orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and to $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$.
@@
    • Feed $\mathcal{A}$ with $\mathbf{p}_{t}^{\top}A_{t}$ and $A_{t}\mathbf{w}_{t}$ (as a reponse to its queries $\mathbf{p}_{t},\mathbf{w}_{t}$), where
      $A_{t}=\sum_{j=1}^{t}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}~{}.$
Note that by construction, $\mathbf{v}_{t}$ is simply a standard Gaussian random vector in $\mathbb{R}^{n}$, scaled by $\beta$ and projected on the subspace orthogonal to $\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$ (in particular, $\mathbf{v}_{t}$ is orthogonal to all these vectors). This is possible by the assumption $2T+1\leq n$, which implies that $M_{t}$ is always a “thin” matrix representing the basis of a strict subspace of $\mathbb{R}^{n}$. Similarly, the assumption $2T+1\leq d$ implies that choosing $\mathbf{u}_{t}\in\mathbb{R}^{d}$ as orthogonal to $12t-$ 1 given vectors (as described above) is indeed possible.
For the construction to be useful, we need to ensure that the oracle responses as defined above (using intermediate matrices $A_{t}$) are all consistent with the same fixed matrix $A$ in hindsight. This is formalized in the following lemma:
@@
Proof.
By the orthogonality assumptions in the construction, for any $t$,
Similarly,
Since $\mathbf{p}_{t},\mathbf{w}_{t}$ at iteration $t$ are determined by the previous oracle calls, it follows by induction that the sequences of vectors $\{\mathbf{p}_{t}\}_{t=1}^{T},\{\mathbf{w}_{t}\}_{t=1}^{T+1}$ are those produced by the algorithm given access to the oracle $\mathcal{O}_{2}^{A}$ for the matrix $A$. Finally, the expression for $A\mathbf{w}_{T+1}$ follows from the definition of $A$, and the fact that $\mathbf{u}_{T+1}$ is chosen to be orthogonal to $\mathbf{w}_{T+1}$. ∎
@@
Proof.
$=~{}\frac{1}{\sqrt{T+1}}\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})+\mathbf{v}_{T}\right)$
$=~{}\frac{1}{\sqrt{T+1}}\cdot\mathbf{v}_{0}~{}=~{}\frac{\alpha}{\sqrt{T+1}}\cdot\mathbf{1}~{}.$
Therefore, $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\frac{\alpha}{\sqrt{T+1}}$ for this choice of unit vector $\mathbf{w}$, so the maximum over all unit vectors can therefore only be larger. ∎
@@
Suppose $A$ is constructed as in Construction 1. Then for any $\delta\in(0,1)$ such that
$2\alpha^{2}+8T\log\left(\frac{2Tn}{\delta}\right)\beta^{2}~{}\leq~{}1~{},$
it holds with probability at least $1-\delta$ that $\max_{l\in[n]}\|A_{l}\|\leq 1$, where $A_{l}$ is the $l$-th row of $A$.
@@
$\sup_{\mathbf{w}\in\mathbb{R}^{d}}~{}\min_{\mathbf{p}\in\Delta^{n-1}}~{}\mathbf{p}^{\top}\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}\leq 0~{},$
@@
The formal proof (which appears in the appendix) is rather involved. At a high level, we use Gaussian concentration properties and an $\epsilon$-net argument, to show that with high probability over the randomized construction, it holds for any $\mathbf{w}$ that the vector $\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ contains an entry which is upper bounded by a certain expression, which we then prove to be non-positive. Therefore, $\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ is non-positive for all $\mathbf{w}$. Combining the proposition with the lemmas above, and choosing the $\alpha,\beta$ parameters appropriately, Thm. 4.2 follows, as detailed in the appendix.
@@ 5 Nash equilibria: simplex/simplex games
As previously discussed, the objective above corresponds to finding a Nash equilibrium in zero-sum bilinear games. In this section we will assume that each entry of $A$ is in $[-1,+1]$, which is the standard normalization for this setting (Rakhlin and Sridharan, 2013; Hadiji et al., 2024).
@@
Suppose the matrix $A$ is constructed as in Construction 1, where instead of $\mathbf{u}_{t}$ being an arbitrary unit vector (orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$), it is chosen specifically as the unit vector $\frac{1}{\|(I-N_{t}N_{t}^{\top})\boldsymbol{\xi}^{\prime}_{t}\|}(I-N_{t}N_{t}^{\top})\boldsymbol{\xi}^{\prime}_{t}$, where the columns of $N_{t}\in\mathbb{R}^{d\times q_{t}}$, $q_{t}\leq 2t-1$ are an orthogonal basis for $\mathbf{u}_{1},\ldots,\mathbf{v}_{t-1},\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$, and $\boldsymbol{\xi}^{\prime}_{t}$ is an independent standard Gaussian random variable in $\mathbb{R}^{d}$.
By construction, $\mathbf{u}_{t}$ still satisfies the desideratum that it is a unit vector orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{v}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$. However, this particular construction gives us probabilistic control over the magnitude of its individual entries. Specifically, we have the following variants of Lemma 4.4 and Lemma 4.5 respectively:
@@
Combining the above two lemmas with Proposition 4.6 (which is unaffected by the additional constraint in Construction 2), and choosing $\alpha,\beta$ accordingly, the proof of Thm. 5.2 readily follows, as detailed in the appendix.
@@ 6 Discussion
Another open problem is extending our lower bounds to randomized algorithms. Since our constructions required simulation of the algorithm’s responses, it is not immediately clear how this can be achieved. We note that previous extensions of oracle complexity lower bounds to randomized algorithms (Woodworth and Srebro, 2017; Carmon et al., 2020a; Arjevani et al., 2023) crucially relied on non-linear modifications of the “hard” target function, which is not possible when we restrict ourselves to bilinear functions.
Appendix A Missing proofs from Sec. 4.2
@@
$\Pr\left(\sum_{l=1}^{n}\mathbf{1}_{(\Xi\mathbf{r})_{l}\geq 1}\leq\frac{n}{20}\right)~{}=~{}\Pr\left(\sum_{l=1}^{n}\mathbf{1}_{(\Xi\mathbf{r})_{l}\geq 1}\leq 0.15n\cdot\frac{1}{3}\right)~{}\leq~{}\exp\left(-\frac{2}{9}\cdot 0.15n\right)~{}=~{}\exp\left(-\frac{n}{30}\right)~{}.$
@@
$\Pr\left(\|M_{j}M_{j}^{\top}\boldsymbol{\xi}_{j}\|\geq z~{}|~{}\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{j-1}\right)\leq 2\exp\left(-\frac{z^{2}}{2\text{Tr}(M_{j}M_{j}^{\top})}\right)=2\exp\left(-\frac{z^{2}}{2\|M_{j}\|_{F}^{2}}\right)\leq 2\exp\left(-\frac{z^{2}}{4j}\right)~{},$
@@
A bit more formally, pick (say) $\delta=1/10$ in both Proposition 4.6 and Lemma 4.5, and choose $\beta=\frac{1}{c\sqrt{T\log(n)}}$ for some sufficiently large universal constant $c$, as well as $\alpha=\frac{\beta}{4\sqrt{T}}=\frac{1}{4cT\sqrt{\log(n)}}$. It is easily verified that with this choice, the assumptions in Proposition 4.6, Lemma 4.4 and Lemma 4.5 are all satisfied, and (with a union bound), the resulting matrix $A$ satisfies both $\max_{l\in[n]}\|A_{l}\|\leq 1$, $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq~{}0$ and $\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{\alpha}{\sqrt{T+1}}~{}\geq\frac{1}{c^{\prime}T\sqrt{T\log(n)}}$ (for some universal constant $c^{\prime}>0$) with some positive probability. Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist.
Appendix B Missing proofs from Sec. 5
@@ Appendix C Auxiliary lemma
$\mathbf{x}$ has a zero-mean Gaussian distribution, with covariance matrix $\beta^{2}(I-MM^{\top})^{2}=\beta^{2}(I-MM^{\top})$. In particular, for any fixed index $l$, the coordinate $x_{l}$ is a zero-mean univariate Gaussian with variance $\beta^{2}\mathbf{e}_{l}^{\top}(I-MM^{\top})\mathbf{e}_{l}=\beta^{2}(1-\|\mathbf{e}_{l}^{\top}M\|^{2})\leq\beta^{2}$. Therefore, by a standard Gaussian tail bound,
$\Pr\left(|x_{l}|\geq z\right)~{}\leq~{}2\exp\left(-\frac{z^{2}}{2\beta^{2}}\right)$
for any $z>0$. Applying a union bound over all indices $l\in[q]$, we get
@@
$\Pr\left(\|(I-MM^{\top})\boldsymbol{\xi}\|^{2}\leq\frac{1}{2}\cdot\frac{q}{2}\right)~{}\leq~{}\Pr\left(\|(I-MM^{\top})\boldsymbol{\xi}\|^{2}\leq\frac{1}{2}\cdot s\right)~{}\leq~{}\exp\left(-\frac{s}{24}\right)~{}\leq~{}\exp\left(-\frac{q}{48}\right)~{}.$

### v2

@@ 4.2 Oracle complexity with a two-sided oracle
We now turn to discuss the proof of Theorem 4.2. We begin by noting that the proof technique of Theorem 4.1 is of no use here, as it is based on revealing the rows of $A$ to the algorithm one at a time. This is not possible with a two-sided oracle, which allows us (for instance) to compute an arbitrary weighted combination of all rows of $A$ using a single query. A second difficulty is that we consider a very simple class of homogeneous bilinear functions, which means that any lower bound necessarily has to be of this form (as opposed to more general min-max optimization problems, where for lower bounds we can use functions with a richer structure). To handle these difficulties, we introduce a different proof technique, which still forces the algorithm to discover information about $A$ in an incremental manner, but in terms of certain vector outer products rather than rows. Specifically, we will utilize the following randomized construction of $A$:
@@
    • Let $\mathbf{u}_{t}$ be some unit vector in $\mathbb{R}^{d}$, which is orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and to $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$.
@@
    • Let $A_{t}=\sum_{j=1}^{t}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}$, and feed $\mathcal{A}$ with $\mathbf{p}_{t}^{\top}A_{t}$ and $A_{t}\mathbf{w}_{t}$ (as a response to its queries $\mathbf{p}_{t},\mathbf{w}_{t}$).
Note that by construction, $\mathbf{v}_{t}$ is a standard Gaussian random vector in $\mathbb{R}^{n}$, scaled by $\beta$ and projected on the subspace orthogonal to $\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$ (hence, $\mathbf{v}_{t}$ is orthogonal to all these vectors). This is possible by the assumption $2T+1\leq n$, which implies that $M_{t}$ is always a “thin” matrix representing the basis of a strict subspace of $\mathbb{R}^{n}$. Similarly, the assumption $2T+1\leq d$ implies that choosing $\mathbf{u}_{t}\in\mathbb{R}^{d}$ as orthogonal to $12t-$ 1 given vectors (as described above) is indeed possible. The choice of a Gaussian distribution is crucially used in order to control the sign and magnitudes of various quantities associated with the matrix, as will be seen later in the proof.
For the construction to be useful, we need to ensure that the oracle responses as defined above (using intermediate matrices $A_{t}$) are all consistent with the same fixed matrix $A$ in hindsight. This is formalized in the following lemma, whose proof easily follows from the orthogonality properties in the construction, as detailed in the appendix.
@@
The proof of Lemma 4.4 notes that $A\mathbf{w}=\frac{\alpha}{\sqrt{T+1}}\mathbf{1}$ for $\mathbf{w}:=\frac{1}{\sqrt{T+1}}\sum_{t=1}^{T+1}\mathbf{u}_{t}$, which is a unit vector following the orthonormality properties in the construction, and hence the maximum is at least this value, as detailed in the appendix.
@@
Suppose $A$ is constructed as in Construction 1. Then for any $\delta\in(0,1)$ such that $2\alpha^{2}+8T\log\left(\frac{2Tn}{\delta}\right)\beta^{2}\leq 1$ it holds with probability at least $1-\delta$ that $\max_{l\in[n]}\|A_{l}\|\leq 1$, where $A_{l}$ is the $l$-th row of $A$.
@@
$\sup_{\mathbf{w}\in\mathbb{R}^{d}}~{}\min_{\mathbf{p}\in\Delta^{n-1}}~{}\mathbf{p}^{\top}\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}~{}\leq~{}0~{},$
@@
The formal proof (which appears in the appendix) is rather involved. At a high level, we use Gaussian concentration properties and an $\epsilon$-net argument, to show that with high probability over the randomized construction, it holds for any $\mathbf{w}$ that the vector $\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ contains an entry which is upper bounded by a certain expression, which we then prove to be non-positive. Therefore, $\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})\mathbf{u}_{j}^{\top}\mathbf{w}$ is non-positive for all $\mathbf{w}$. Combining the proposition with the lemmas above, and choosing the $\alpha,\beta$ parameters appropriately, Theorem 4.2 follows, as detailed in the appendix.
@@ 5 Nash equilibria: simplex/simplex games
As previously discussed, the objective above corresponds to finding a Nash equilibrium in zero-sum bilinear games. In this section we will assume that each entry of $A$ is in $[-1,+1]$, which is the standard normalization for this setting (cf. Nemirovski, 2004; Hadiji et al., 2024).
@@
Suppose the matrix $A$ is constructed as in Construction 1, where instead of $\mathbf{u}_{t}$ being an arbitrary unit vector (orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$), it is chosen specifically as the unit vector $\frac{1}{\|(I-N_{t}N_{t}^{\top})\boldsymbol{\xi}^{\prime}_{t}\|}(I-N_{t}N_{t}^{\top})\boldsymbol{\xi}^{\prime}_{t}$, where the columns of $N_{t}\in\mathbb{R}^{d\times q_{t}}$, $q_{t}\leq 2t-1$ are an orthogonal basis for $\mathbf{u}_{1},\ldots,\mathbf{v}_{t-1},\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$, and $\boldsymbol{\xi}^{\prime}_{t}\sim\mathcal{N}(\mathbf{0},I_{d})$ is an independent Gaussian.
By construction, $\mathbf{u}_{t}$ still satisfies the desideratum that it is a unit vector orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{v}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$. However, this particular construction gives us probabilistic control over the magnitude of its individual entries. Specifically, we have the following:
@@
Note that the lemmas above are variants of Lemma 4.4 and Lemma 4.5 respectively. Combining them with Proposition 4.6 (which is unaffected by the additional constraint in Construction 2), and choosing $\alpha,\beta$ accordingly, the proof of Theorem 5.2 readily follows, as detailed in the appendix.
@@ 6 Discussion
Another open problem is extending our lower bounds to randomized algorithms. Since our constructions required simulation of the algorithm’s responses, it is not immediately clear how this can be achieved. We note that previous extensions of oracle complexity lower bounds to randomized algorithms (Nemirovski and Yudin, 1983; Carmon et al., 2020a; Arjevani et al., 2023) crucially relied on non-linear modifications of the “hard” target function, which is not possible when we restrict ourselves to bilinear functions.
Acknowledgments
This research is supported in part by European Research Council (ERC) grant 754705. GK is supported by an Azrieli Foundation graduate fellowship.
Appendix A Missing proofs from Section 4.2
@@ Acknowledgments
By the orthogonality assumptions in the construction, for any $t$,
Similarly,
Since $\mathbf{p}_{t},\mathbf{w}_{t}$ at iteration $t$ are determined by the previous oracle calls, it follows by induction that the sequences of vectors $\{\mathbf{p}_{t}\}_{t=1}^{T},\{\mathbf{w}_{t}\}_{t=1}^{T+1}$ are those produced by the algorithm given access to the oracle $\mathcal{O}_{2}^{A}$ for the matrix $A$. Finally, the expression for $A\mathbf{w}_{T+1}$ follows from the definition of $A$, and the fact that $\mathbf{u}_{T+1}$ is chosen to be orthogonal to $\mathbf{w}_{T+1}$.
A.2 Proof of Lemma 4.4
$=~{}\frac{1}{\sqrt{T+1}}\left(\sum_{j=1}^{T}(\mathbf{v}_{j-1}-\mathbf{v}_{j})+\mathbf{v}_{T}\right)=~{}\frac{1}{\sqrt{T+1}}\cdot\mathbf{v}_{0}~{}=~{}\frac{\alpha}{\sqrt{T+1}}\cdot\mathbf{1}~{}.$
Therefore, $\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\frac{\alpha}{\sqrt{T+1}}$ for this choice of unit vector $\mathbf{w}$, so the maximum over all unit vectors can therefore only be larger.
A.3 Proof of Lemma 4.5
@@
$\Pr\left(\sum_{l=1}^{n}\mathbf{1}_{(\Xi\mathbf{r})_{l}\geq 1}\leq\frac{n}{20}\right)$ $=\Pr\left(\sum_{l=1}^{n}\mathbf{1}_{(\Xi\mathbf{r})_{l}\geq 1}\leq 0.15n\cdot\frac{1}{3}\right)~{}$
$\leq\exp\left(-\frac{2}{9}\cdot 0.15n\right)=\exp\left(-\frac{n}{30}\right)~{}.$
@@
$\Pr\left(\|M_{j}M_{j}^{\top}\boldsymbol{\xi}_{j}\|\geq z~{}|~{}\boldsymbol{\xi}_{1},\ldots,\boldsymbol{\xi}_{j-1}\right)$ $\leq 2\exp\left(-\frac{z^{2}}{2\text{Tr}(M_{j}M_{j}^{\top})}\right)=2\exp\left(-\frac{z^{2}}{2\|M_{j}\|_{F}^{2}}\right)$
$\leq 2\exp\left(-\frac{z^{2}}{4j}\right)~{},$
@@
A bit more formally, pick (say) $\delta=1/10$ in both Proposition 4.6 and Lemma 4.5, and choose $\beta=\frac{1}{c\sqrt{T\log(n)}}$ for some sufficiently large universal constant $c$, as well as $\alpha=\frac{\beta}{4\sqrt{T}}=\frac{1}{4cT\sqrt{\log(n)}}$. It is easily verified that with this choice, the assumptions in Proposition 4.6, Lemma 4.4 and Lemma 4.5 are all satisfied, and (with a union bound), the resulting matrix $A$ satisfies both
$\max_{l\in[n]}\|A_{l}\|\leq 1~{},~{}~{}~{}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}_{T+1}\leq 0~{},\text{~{}~{}~{}and~{}~{}~{}}\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{\alpha}{\sqrt{T+1}}~{}\geq\frac{1}{c^{\prime}T\sqrt{T\log(n)}}$
(for some universal constant $c^{\prime}>0$) with some positive probability. Hence, by the probabilistic method, a suitable fixed matrix $A$ satisfying all of the above must exist.
Appendix B Missing proofs from Section 5
@@ Appendix C Auxiliary lemma
$\mathbf{x}$ has a zero-mean Gaussian distribution, with covariance matrix $\beta^{2}(I-MM^{\top})^{2}=\beta^{2}(I-MM^{\top})$. In particular, for any fixed index $l$, the coordinate $x_{l}$ is a zero-mean univariate Gaussian with variance $\beta^{2}\mathbf{e}_{l}^{\top}(I-MM^{\top})\mathbf{e}_{l}=\beta^{2}(1-\|\mathbf{e}_{l}^{\top}M\|^{2})\leq\beta^{2}$. Therefore, by a standard Gaussian tail bound, $\Pr\left(|x_{l}|\geq z\right)\leq 2\exp\left(-{z^{2}}/{2\beta^{2}}\right)$ for any $z>0$. Applying a union bound over all indices $l\in[q]$, we get
@@
$\Pr\left(\|(I-MM^{\top})\boldsymbol{\xi}\|^{2}\leq\frac{1}{2}\cdot\frac{q}{2}\right)$ $\leq\Pr\left(\|(I-MM^{\top})\boldsymbol{\xi}\|^{2}\leq\frac{1}{2}\cdot s\right)\leq\exp\left(-\frac{s}{24}\right)$
$\leq\exp\left(-\frac{q}{48}\right)~{}.$